If you deposit $10,000 into a financial institution today with a
5% annual interest rate, but interest is compounded continuously.
How much would you have accumulated at the end of 3 years?
Answer #1

If an amount P is deposited in an account for t years at an interest rate of r% which is continuously compounding, then its value at the end of t years is given by-

A = P*e ^(rt)

In this question, P= 10000, r = 0.05 ,t= 3

So, A = 10000*e^0.15 = 10000(1.161834) = 11618.34